OpenStudy (anonymous):

@timaashorty @Jack1 @mahnoorkeen2001 x = 9/4x = 3 quadratic question

OpenStudy (anonymous):

@sarah1017

OpenStudy (anonymous):

@erielle

OpenStudy (anonymous):

@hanner_B_nanner @hartnn

hartnn (hartnn):

what the question ? verify it again...

OpenStudy (anonymous):

oops x + 9/4x = 3

OpenStudy (anonymous):

keyboard lagging

hartnn (hartnn):

yeah, first bring it in the form of \(ax^2+bx+c = 0\) multiply each term by 4x what do u get ?

OpenStudy (anonymous):

even 9 ?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes, both sides, every term!

OpenStudy (anonymous):

so 3 will be 12 or 12x

hartnn (hartnn):

3 \times 4x = 12 x :)

OpenStudy (anonymous):

so 4x² + 36x = 12x

hartnn (hartnn):

the 2nd term is 9/4x !! right ? so, 9/4x * 4x is just 9

OpenStudy (anonymous):

can u write th e equation pleae im confussed

OpenStudy (anonymous):

its suppose to be 4x² - 12x + 9 = 0 ????@hartnn

hartnn (hartnn):

yes!

OpenStudy (anonymous):

now i am stuck im suppose to solve by factoring

hartnn (hartnn):

what are 2 numbers whose sum = -12 and product = 4*9 = 36 ?

OpenStudy (anonymous):

im stuck there itself

OpenStudy (anonymous):

is it - 9 and -4 ???? @hartnn

hartnn (hartnn):

what about -6 and -6 ? :P

OpenStudy (anonymous):

lol was gonna type that thanks for helping

hartnn (hartnn):

woah ! thise are not final answer, they will just help you factor the quadratic

hartnn (hartnn):

*thise

OpenStudy (anonymous):

i will do the rest

hartnn (hartnn):

good! :)

OpenStudy (anonymous):

so 2x(2x-3) - 3(2x-3) = 0

hartnn (hartnn):

yep :)

OpenStudy (anonymous):

so (x = 3/2 , 3/2)

hartnn (hartnn):

thats correct :)

OpenStudy (anonymous):

i hv more questions

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

(x - 4)² = 25 @hartnn

hartnn (hartnn):

take square root on both sides :)

OpenStudy (anonymous):

how do so

hartnn (hartnn):

\(\large a^2 = b \implies a= \pm \sqrt b \\ \large a = \sqrt b, \quad a = - \sqrt b \)

hartnn (hartnn):

here, a = x-4 b= 25

OpenStudy (anonymous):

wait not getting you

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

what you are not getting ? i just took square root on both sides... \(\sqrt {a^2} = a\)

OpenStudy (anonymous):

so is b = 5 @hartnn

hartnn (hartnn):

that was just a general form...